OpenStudy (anonymous):

how do solve 5(x+4)=2(x+x+2)+42

OpenStudy (anonymous):

Distribute the 5 and the 2, so 5(x+4)=2(x+x+2)+42 becomes 5x+20=2x+2x+4+42, simplify, so.... 5x+20=4x+46 then subtract one x from the other to isolate x.... x+20=46 then get x on one side so... x=26